### v1

# Con4m: Context-aware Consistency Learning Framework for Segmented Time Series Classification

Abstract
Time Series Classification (TSC) encompasses two settings: classifying entire sequences or classifying segmented subsequences. The raw time series for segmented TSC usually contain Multiple classes with Varying Duration of each class (MVD). Therefore, the characteristics of MVD pose unique challenges for segmented TSC, yet have been largely overlooked by existing works. Specifically, there exists a natural temporal dependency between consecutive instances (segments) to be classified within MVD. However, mainstream TSC models rely on the assumption of independent and identically distributed (i.i.d.), focusing on independently modeling each segment. Additionally, annotators with varying expertise may provide inconsistent boundary labels, leading to unstable performance of noise-free TSC models. To address these challenges, we first formally demonstrate that valuable contextual information enhances the discriminative power of classification instances. Leveraging the contextual priors of MVD at both the data and label levels, we propose a novel consistency learning framework Con4m, which effectively utilizes contextual information more conducive to discriminating consecutive segments in segmented TSC tasks, while harmonizing inconsistent boundary labels for training. Extensive experiments across multiple datasets validate the effectiveness of Con4m in handling segmented TSC tasks on MVD.
footnotetext: $\dagger$ Corresponding author.
1 Introduction
Figure 1: (a) In the healthcare domain, different physicians may reach a consensus on the seizure onset, but they have varying annotations regarding the end times of seizure waves. (b) Reasonable model predictions exhibit coherence across consecutive segments rather than repeated interruptions. (c) Based on the proximity to the boundary, we divide each class sequence into 5 levels, from which an equal number of segments are sampled. A one-layer MLP is trained on the segments from each level respectively for the same number of epochs. (d) We visualize the predicted probability of the trained MLP for each level. We observe that as the segments approach the boundaries, the model finds it increasingly challenging to make correct classifications, resulting in more extreme wrong predictions. This strongly underscores the significance of handling boundary segments.
Time Series Classification (TSC) is one of the most challenging problems in the field of machine learning. TSC aims to assign labels to a series of temporally ordered data points. These points either form a complete sequence or are subsequences (segments) resulting from the segmentation of a long time series. Existing works [46, 18] largely focus on the assumption of independent and identically distributed (i.i.d.), in which case each sequence or segment is regarded as an independent instance to be classified, not differentiating between these two settings. In fact, for many practical applications, the raw time series before segmentation for segmented TSC tasks contain Multiple classes with Varying Duration of each class (MVD). For example, in the healthcare domain, the brain signals of epileptic patients often record over several days, encompassing multiple seizure onsets, each with varying durations and intervals. In the field of activity recognition, sensors continuously record users’ behavior data, including walking, riding, and running, among other activities, each with varying durations. Therefore, the characteristics of the raw MVD lead to the uniqueness of segmented TSC tasks. Given this, our research concentrates on effectively modeling segmented TSC tasks based on MVD, presenting distinctive challenges.
(1) Leveraging contextual information. In contrast to TSC tasks for complete sequences, in which classified sequences are relatively independent, there exist natural temporal dependencies between consecutive classified segments for segmented TSC. We take the seizure detection task as an example, in which given the brain signals of epileptic patients, the model should identify whether a segment includes seizure waves or not. As illustrated in Figure 1(b), given the sustained nature of seizure onsets, the model predictions for consecutive segments should exhibit coherence, with seizure and normal predictions appearing in continuous and concentrated patterns. However, mainstream TSC models [54, 67, 57] focus on the i.i.d. assumption and model the internal context within each segment to be classified, largely overlooking the dependencies between consecutive segments.
In the domain of video analysis, works in temporal action segmentation (TAS) [16] have modeled the temporal dependency between different video frames and made frame-wise predictions. However, unlike the I3D features [7] used as input in these works, time series lack a unified pretrained model for feature extraction, and the dependency between segments is more variable and ambiguous. Furthermore, TAS works focus on modeling the dependency of instances from a data perspective, without explicitly leveraging contextual label information. Therefore, how to leverage contextual information in segmented TSC tasks to make more reasonable classifications is crucial and challenging.
(2) Inconsistent boundary labels. In domains with precious labels, different annotators collaboratively contribute annotations. The raw annotations of MVD typically include information about the start and end times for each class. However, due to inherent ambiguity and a lack of unified quantification standards, differences in annotator experience may lead to inconsistent boundary labels. Returning to Figure 1(a), in the seizure detection task, owing to the natural fuzzy transition from seizure onset to completely normal, different physicians have varying experiences regarding when seizure waves terminate.
Furthermore, inconsistent boundary labeling causes that boundary segments with similar patterns may have opposite labels, leading to unstable model training. To validate the detrimental impact, as shown in Figure 1(c), we divide each class sequence in the seizure detection task into 5 levels, where higher levels indicate proximity to the boundary. We then sample an equal number of balanced binary segments for each level. Subsequently, a one-layer MLP is trained for the same number of epochs on the segments from each level respectively. Figure 1(d) visualizes the results after training. We observe that as the level increases (closer to the boundary), the model’s accuracy steadily decreases, and erroneous predictions become more extreme. The results highlight the significant detrimental impact of inconsistent boundary labels on noise-free model performance.
Noisy label learning (NLL) [52] aims to learn robust models from data containing corrupted labels. While the inconsistent labels in MVD are not intentionally corrupted but rather stem from implicit discrepancies due to experiential differences, NLL remains the most relevant approach to address such discrepancies. To the best of our knowledge, Scale-teaching [43] and SREA [8] are the only NLL works specifically designed for time series and are thus the most relevant to our work. However, they also face the issue of overlooking contextual dependencies across consecutive time segments, posing the challenge of handling inconsistent boundary labels using context during training.
To overcome the challenges above, we propose Con4m–a label Consistency learning framework, which leverages effective Contextual information, achieving Coherent predictions and Continuous representations for segmented TSC tasks, while harmonizing inconsistent boundary labels for training. Specifically, we first formally demonstrate that valuable contextual information enhances the discriminative power of classification instances. Based on the insights, by incorporating prior knowledge of data locality and label coherence, we guide and constrain the model to focus on contextual information more conducive to discriminating consecutive segments in segmented TSC tasks. Meanwhile, leveraging model predictions that thoroughly encompass contextual information, Con4m progressively changes the training labels in an adaptive manner to harmonize inconsistent labels across consecutive segments. This leads to a more robust model. Our contributions are summarized as follows:
(1) We are the first to propose a practical consistency learning framework Con4m for the segmented TSC based on the raw MVD. (2) By comprehensively integrating prior knowledge from both the data and label perspectives, we guide the model to focus on effective contextual information. Based on context-aware predictions, a progressive harmonization approach for handling inconsistent training labels is designed to yield a more robust model. (3) Extensive experiments on two public and one private MVD datasets demonstrate the superior performance of Con4m. The Con4m’s ability to harmonize inconsistent labels is further verified by the label substitution experiment and a case study.
2 Theoretical Analysis
In this section, we aim to formally demonstrate the benefit of contextual information for classification tasks, and to establish the existence of an upper bound for this benefit. Consequently, by introducing prior knowledge, we can guide the model to focus on valuable contextual information more conducive to improving the benefit for segmented TSC tasks.
Assuming that the random variables of the instances to be classified and the corresponding labels are denoted as $\textnormal{x}_{t}$ and $\textnormal{y}_{t}$. $\mathbb{A}_{t}$ represents the contextual instance set introduced for $\textnormal{x}_{t}$. $x_{\mathbb{A}_{t}}$ denotes the random variable for the contextual instance set. Mutual information measures the correlation between two random variables. In a classification task, a higher correlation between instances and labels indicates that the instances are more easily distinguishable by the labels. This benefits the classification task, making it more readily addressable. Therefore, from an information-theoretic perspective, we elucidate the benefit of contextual information through the following theorem.
Theorem 2.1.
The more the introduced contextual instance set enhance the discriminative power of the target instance, the greater the benefit for the classification task.
Proof.
Firstly, we establish that the introduction of contextual information does not compromise classification tasks, i.e., it does not diminish the correlation between instances and labels.
$\mathbb{I}(\textnormal{y}_{t};\textnormal{x}_{t},\textnormal{x}_{\mathbb{A}_{t}})=\mathbb{I}(\textnormal{y}_{t};\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})+\mathbb{I}(\textnormal{y}_{t};\textnormal{x}_{t})\geq\mathbb{I}(\textnormal{y}_{t};\textnormal{x}_{t}).$ (1)
The inequality holds due to the non-negativity of conditional mutual information.
According to (1), the increase in $\mathbb{I}(\textnormal{y}_{t};\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})$ determines the extent to which the introduction of contextual information can be beneficial for classification tasks. Expanding $\mathbb{I}(\textnormal{y}_{t};\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})$, we have:
$\mathbb{I}(\textnormal{y}_{t};\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})$ $=\sum_{\textnormal{x}_{t}}p(\textnormal{x}_{t})\sum_{\textnormal{x}_{\mathbb{A}_{t}}}\sum_{\textnormal{y}_{t}}p(\textnormal{y}_{t},\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})\log{\frac{p(\textnormal{y}_{t},\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})}{p(\textnormal{y}_{t}|\textnormal{x}_{t})p(\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})}}$
$=\sum_{\textnormal{x}_{t}}p(\textnormal{x}_{t})\sum_{\textnormal{x}_{\mathbb{A}_{t}}}\sum_{\textnormal{y}_{t}}p(\textnormal{y}_{t}|\textnormal{x}_{t},\textnormal{x}_{\mathbb{A}_{t}})p(\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})\log{\frac{p(\textnormal{y}_{t}|\textnormal{x}_{t},\textnormal{x}_{\mathbb{A}_{t}})}{p(\textnormal{y}_{t}|\textnormal{x}_{t})}}$
$=\sum_{\textnormal{x}_{t}}p(\textnormal{x}_{t})\sum_{\textnormal{x}_{\mathbb{A}_{t}}}p(\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})D_{\mathrm{KL}}(p(\textnormal{y}_{t}|\textnormal{x}_{t},\textnormal{x}_{\mathbb{A}_{t}})\|p(\textnormal{y}_{t}|\textnormal{x}_{t})).$
Given a fixed instance $\textnormal{x}_{t}$ and the inherent distribution $p(\textnormal{y}_{t}|\textnormal{x}_{t})$ of the data, the KL divergence is a convex function for $\textnormal{x}_{\mathbb{A}_{t}}$ that attains its minimum at $p(\textnormal{y}_{t}|\textnormal{x}_{t},\textnormal{x}_{\mathbb{A}_{t}})=p(\textnormal{y}_{t}|\textnormal{x}_{t})$. As $p(\textnormal{y}_{t}|\textnormal{x}_{t},\textnormal{x}_{\mathbb{A}_{t}})$ approaches the boundary of the probability space, where the predictive probability of one class approaches $1$ and the rest approach 00, the value of KL divergence increases. A stronger discriminative power regarding $\textnormal{x}_{t}$ implies less uncertainty regarding $\textnormal{y}_{t}$, which is equivalent to approaching the boundary of the probability space.
Due to the convexity of the KL divergence and the boundedness of $p(\textnormal{y}_{t}|\textnormal{x}_{t},\textnormal{x}_{\mathbb{A}_{t}})$, there exists a contextual instance set in the data that maximizes $D_{\mathrm{KL}}(p(\textnormal{y}_{t}|\textnormal{x}_{t},\textnormal{x}_{\mathbb{A}_{t}})\|p(\textnormal{y}_{t}|\textnormal{x}_{t}))$. We denote the instance set as $\mathbb{A}_{t}^{*}$ and the maximum value of KL divergence as $D_{t}^{*}$. Besides, we note that $\sum_{\textnormal{x}_{\mathbb{A}_{t}}}p(\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})=1$. Hence, we can obtain the upper bound for the information gain $\mathbb{I}(\textnormal{y}_{t};\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})\leq\sum_{\textnormal{x}_{t}}p(\textnormal{x}_{t})\sum_{\textnormal{x}_{\mathbb{A}_{t}}}p(\textnormal{x}_{\mathbb{A}_{t}}|\textnormal{x}_{t})D_{t}^{*}\leq\sum_{\textnormal{x}_{t}}p(\textnormal{x}_{t})D_{t}^{*}$. The convexity of the KL divergence also implies monotonicity, indicating that as $A_{t}$ approaches $A^{*}_{t}$, the KL divergence increases, leading to a greater information gain for the classification task. ∎
According to Theorem 2.1, valuable contextual information enhances the discriminative power of the instances. While the optimal instance set $A^{*}_{t}$ is challenging to directly obtain or optimize, focusing the model on contextual instances more likely to be included in $A^{*}_{t}$ is beneficial for enhancing the performance of the classification task. Furthermore, $\textnormal{x}_{\mathbb{A}_{t}}$ not only contains information at the data level but also encompasses information at the label level (which can be replaced with $\textnormal{y}_{\mathbb{A}_{t}}$). Therefore, we can guide the model to focus on contextual information more conducive to segmented TSC tasks by simultaneously introducing prior knowledge from both the data and label perspectives.
3 The Con4m Method
In this section, we introduce the details of Con4m. Based on the insights of Theorem 2.1, we introduce contextual prior knowledge of data locality (Sec. 3.1) and label coherence (Sec. 3.2) to guide the model to focus on contextual information more conducive to discriminating consecutive segments in segmented TSC tasks. In Sec. 3.3, inspired by the idea of noisy label learning, we propose a label harmonization framework to achieve a more robust model. Before delving into the details of Con4m, we provide the formal definition of the segmented TSC task in our work.
Definition 3.1.
Given a time interval comprising of $T$ consecutive time points and labels, denoted as $(X,Y)=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\dots,(X_{T},Y_{T})\}$, a $w$-length sliding window with stride length $r$ is employed for segmentation. $(X,Y)$ is partitioned into $L$ time segments, represented as $(x,y)=\{(x_{i},y_{i})=(\{X_{(i-1)\times r+1},\dots,X_{(i-1)\times r+w}\},\text{Majority}(\{Y_{(i-1)\times r+1},\dots,Y_{(i-1)\times r+w}\}))|i=1,\dots,L\}$. The model is tasked with predicting segmented labels $y_{i}$ for each time segment $x_{i}$.
3.1 Continuous Contextual Representation Encoder
Local continuity is an inherent attribute of MVD, meaning each class should be locally continuous and only change at its actual boundary. Smoothing with a Gaussian kernel [17, 15, 62] promotes the continuity of representations of time segments in a local temporal window. This not only helps the model make similar predictions of consecutive segments within the same class but also aligns with the gradual nature of class transitions. Furthermore, for graph neural networks based on the homophily assumption, aggregating neighbor information belonging to the same class can improve the discriminative power of the target instance [45, 69]. Therefore, we introduce the Gaussian prior to guide the model to focus on contextual instances $\mathbb{A}_{t}$ proximate to the target instance.
Vanilla self-attention [14] with point-wise attention computations often fail to obtain continuous representations after aggregation. Therefore, we use the Gaussian kernel $\Phi(x,y|\sigma)$ as prior weights to aggregate neighbors to obtain smoother representations. Since the neighbors of boundary segments may belong to different classes, we allow each segment to learn its own scale parameter $\sigma$. Formally, as Figure 2 shows, the two-branch Con-Attention in the $l$-th layer is:
$Q,K,V_{s},V_{g},\sigma=$ $c^{l-1}W_{Q}^{l},c^{l-1}W_{K}^{l},c^{l-1}W_{V_{s}}^{l},c^{l-1}W_{V_{g}}^{l},c^{l-1}W_{\sigma}^{l},$
$S^{l}=\text{SoftMax}\left(\frac{QK^{\top}}{\sqrt{d}}\right),\quad G^{l}$ $=\text{Rescale}\left(\left[\frac{1}{\sqrt{2\pi}\sigma_{i}}\text{exp}\left(-\frac{|j-i|^{2}}{2\sigma_{i}^{2}}\right)\right]_{i,j\in\{1,\dots,L\}}\right),$
$z_{s}^{l}=S^{l}V_{s},\quad$ $z_{g}^{l}=G^{l}V_{g},\quad z^{l}=\text{Fusion}(z_{s}^{l},z_{g}^{l}),$
where $L$ is the number of consecutive segments, $d$ is the dimension of hidden representations, $c^{l-1}\in\mathbb{R}^{L\times d}$ is the output representations of the $l-1$-th layer, and $W_{*}^{l}\in\mathbb{R}^{d\times d}$ are all learnable matrices. $\text{Rescale}(\cdot)$ refers to row normalization by index $i$. To distinguish between two computational branches, we use $g/G$ to represent the branch based on Gaussian prior, and $s/S$ to represent the branch based on self-attention. $S^{l}$ and $G^{l}$ are the aggregation weights. We use the conventional attention mechanism [4] to adaptively fuse $z_{s}^{l}$ and $z_{g}^{l}$. Finally, as illustrated in Figure 2, by stacking the multi-head version of Con-Attention layers, we construct Con-Transformer, which serves as the backbone of the continuous encoder of Con4m to obtain final representations $c$.
Figure 2: Overview of the encoder of Con4m. The leftmost part shows the details of Con-Attention. The right part of the figure shows the architecture of Con-Transformer and the whole encoder of Con4m.
3.2 Context-aware Coherent Class Prediction
In the segmented TSC task of MVD, consecutive time segments not only provide contextual information at the data level but also possess their own class information. As depicted in Figure 1(a), considering the persistence of each class and the gradual nature of class transitions, the model’s predictions should exhibit more coherence and concentration, rather than being interspersed. Therefore, we integrate and constrain the model’s predictions from both the individual and holistic perspectives to achieve more coherent predictions.
Neighbor Class Consistency Discrimination. In graphs, label propagation algorithms [26, 29] are often utilized to refine and smooth the predictions of neighbor instances, thereby enhancing their discrimination. Drawing inspiration from this, by weightedly aggregating predictions from similar time segments, the model can focus on contexts $\mathbb{A}_{t}$ more likely to belong to the same class as the target segment. Although there is no explicit graph structure between time segments, we can train a discriminator to determine whether two segments belong to the same class. The model then aggregates the contextual class predictions based on the discriminator’s outputs, thus making more robust predictions. As the left part of Figure 3 shows, we formalize this process as follows:
$\hat{R}=\text{SoftMax}\left(\left[\text{MLP}_{2}\left(c_{i}\|c_{j}\right)\right]_{i,j\in\{1,\dots,L\}}\right),\quad\hat{p}=\text{SoftMax}\left(\text{MLP}_{1}\left(c\right)\right),\quad\tilde{p}=\hat{R}_{:,:,1}\hat{p},$
[…]
Prediction Behavior Constraint. Unlike graphs, there exists a holistic temporal relationship between consecutive time segments. Therefore, we should further constrain the overall predictive behavior along the time axis. For MVD, as Figure 1(a) shows, within a suitably chosen time interval, consecutive segments almost span at most two classes. Therefore, we ensure the monotonicity of predictions across consecutive segments through hard constraints, thereby utilizing contextual label information $\textnormal{y}_{\mathbb{A}_{t}}$ to integrate and refine predictions across these segments.
As shown in the middle part of Figure 3, for each class in the predictions, there are only four prediction behaviors for consecutive segments, namely high confidence, low confidence, confidence decreasing, and confidence increasing. To constrain the behavior, we use function fitting to integrate $\tilde{p}$. Considering the wide applicability, we opt for the hyperbolic tangent function (i.e., Tanh) as our basis. Formally, we introduce four tunable parameters to exactly fit the monotonicity as:
$\bar{p}=\text{Tanh}(x|a,k,b,h)=a\times\text{Tanh}\left(k\times\left(x+b\right)\right)+h,$
where parameter $a$ constrains the range of the function’s values, $k$ controls the slope of the transition of the function, $b$ and $h$ adjust the symmetry center of the function, and $x$ is the given free vector in the x-coordinate. We use the MSE loss to fit the contextual predictions $\tilde{p}$ as $\ell_{3}=\|\text{Tanh}(x|a,k,b,h)-\tilde{p}\|^{2}$. It deserves to be emphasized that $\tilde{p}$ in the process has no gradient and therefore does not affect the parameters of the encoder. Please see Appendix B for more fitting details.
After function fitting, we obtain independent predictions $\hat{p}$ for each segment and constrained predictions $\bar{p}$ that leverage contextual label information. For the inference stage, we use the average of them as the final coherent predictions, i.e., $\hat{y}=\arg\max{(\hat{p}+\bar{p})/2}$.
Figure 3: Overview of context-aware coherent class prediction and consistent label training framework in Con4m. The left part describes the neighbor class consistency discrimination task and the prediction behavior constraint. The rightmost part presents the training and inference details for label harmonization.
3.3 Label Consistency Training Framework
Due to inherent ambiguity, the annotation of MVD often lacks quantitative criteria, resulting in experiential differences across individuals. Such discrepancies are detrimental to models and we propose a training framework to enable Con4m to adaptively harmonize inconsistent labels.
Learning from easy to hard. We are based on the fact that although people may have differences in the fuzzy transitions between classes, they tend to reach an agreement on the most significant core part of each class. In other words, the empirical differences become more apparent when approaching the transitions. Therefore, we adopt curriculum learning techniques to help the model learn instances from the easy (core) to the hard (transition) part. Formally (see the diagram in Figure 1(b)), for a continuous $K$-length class, we divide it into $N_{l}=5$ equally sized levels as follows:
$\left(\lceil(N_{l}-1)\frac{K}{2N_{l}}\rceil,\lfloor(N_{l}+1)\frac{K}{2N_{l}}\rfloor\right);\cdots;\left[1,\lceil\frac{K}{2N_{l}}\rceil\right)\bigcup\left(\lfloor(2N_{l}-1)\frac{K}{2N_{l}}\rfloor,K\right].$ (2)
Then we sample the same number of time intervals from each level. The higher the level, the more apparent the inconsistency. Therefore, as the right part of Figure 3 shows, during the training stage, Con4m learns the time intervals in order from low to high levels, with a lag gap of $E_{g}=5$ epochs.
Harmonizing inconsistent labels. Inspired by the idea of noisy label learning, we gradually change the raw labels to harmonize the inconsistency. The model preferentially changes the labels of the core segments that are easier to reach a consensus, which can avoid overfitting of uncertain labels. Moreover, the model will consider both the independent and constrained predictions to robustly change inconsistent labels. Specifically, given the initial label $y_{0}$, we update the labels $y_{e}=\arg\max{p_{e}}$ for the $e$-th epoch, where $p_{e}$ is obtained as follows:
$\hat{p}_{e}^{5}=\omega_{e}\cdot$ $\left[\hat{p}_{e-m}\right]_{m\in\{0,\dots,4\}},\quad\bar{p}_{e}^{5}=\omega_{e}\cdot\left[\bar{p}_{e-m}\right]_{m\in\{0,\dots,4\}},$
$p_{e}$ $=\left(1-\eta\right)y_{0}+\eta\left(\left(1-\frac{\eta}{2}\right)\hat{p}_{e}^{5}+\frac{\eta}{2}\bar{p}_{e}^{5}\right),$
where $\omega_{e}=\text{Rescale}([\text{exp}((e-m)/2)]_{m\in\{0,\dots,4\}})$ is the exponentially averaged weight vector to aggregate the predictions of the latest $5$ epochs to achieve a more robust label update. $\hat{p}_{e-m}$ and $\bar{p}_{e-m}$ are the independent and constrained predictions in the $e-m$-th epoch respectively and $\cdot$ denotes the dot product. The dynamic weighting factor, $\eta$, is used to adjust the degree of label update. As Figure 3 shows, $\eta$ linearly increases from 00 to $1$ with $E_{\eta}$ epochs, gradually weakening the influence of the original labels. Besides, in the initial training stage, the model tends to improve independent predictions. As the accuracy of independent predictions increases, the model assigns a greater weight to the constrained predictions. See the hyperparameter analysis for $E_{\eta}$ in Appendix C.
4 Experiment
4.1 Experimental Setup
Datasets. In this work, we use two public [27, 33] and one private MVD data to measure the performance of models. Specifically, the Tufts fNIRS to Mental Workload [27] data (fNIRS) contains brain activity recordings from adult humans performing controlled cognitive workload tasks. The SleepEDF [33] data (Sleep) contains PolySomnoGraphic sleep records for subjects over a whole night. The private SEEG data records brain signals indicative of suspected pathological tissue within the brain of epileptic patients. More detailed descriptions can be found in Table 1 and Appendix D.
Table 1: Overview of MVD datasets used in this work.
Data Sample # of # of Subjects Groups Cross Total Interval Window Slide Total
Frequency Features Classes Validation Intervals Length Length Length Segments
fNIRS 5.2Hz 8 2 68 4 12 4,080 38.46s 4.81s 0.96s 146,880
Sleep 100Hz 2 5 154 3 6 6,000 40s 2.5s 1.25s 186,000
SEEG 250Hz 1 2 8 4 3 8,000 16s 1s 0.5s 248,000
Label disturbance. We introduce a novel disturbance method to the raw labels $Y$ of the public datasets to simulate scenarios where labels are inconsistent. Specifically, we first look for the boundary points between different classes in a complete long MVD data. Then, we randomly determine with a $0.5$ probability whether each boundary point should move forward or backward. Finally, we randomly select a new boundary point position from $r\%$ of the length of the class in the direction of the boundary movement. In this way, we can interfere with the boundaries and simulate label inconsistency. Meanwhile, a larger value of $r\%$ indicates a higher degree of label inconsistency. For SEEG dataset, inconsistent labels already exist in the raw data and we do not disturb it.
Baselines. We compare Con4m with state-of-art models from various domains, including two noisy label learning (NLL) models for time series classification (TSC): SREA [8] and Scale-teaching [43] (Scale-T), three image classification models with noisy labels: SIGUA [24], UNICON [32] and Sel-CL [39], three supervised TSC models: MiniRocket [12], TimesNet [60] and PatchTST [47], and three temporal action segmentation (TAS) models: MS-TCN2 [38], ASFormer [64] and DiffAct [41]. See more detailed descriptions of the baselines in Appendix E.
Implementation details. We use cross-validation [35] to evaluate the model’s generalization ability by partitioning the subjects in the data into non-overlapping subsets for training and testing. As shown in Table 1, for fNIRS and SEEG, we divide the subjects into $4$ groups and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct experiments. We divide the Sleep dataset into $3$ groups and follow the $1$-$1$-$1$ experimental setting. Therefore, we only report the mean values of cross-validation results in the main context. See more details and the full results in Appendix G.
4.2 Label Disturbance Experiment
The average results over all cross-validation experiments are presented in Table 2. Overall, Con4m outperforms almost all baselines across all datasets and all disturbance ratios.
Table 2: Comparison with baseline methods in the testing $F_{1}$ score (%) on three datasets. The best results are in bold and we underline the second best results. The worst results are denoted in italics.
Model $r\%$ fNIRS [27] Sleep [33] SEEG
$0\%$ $20\%$ $40\%$ $0\%$ $20\%$ $40\%$ raw
TAS MS-TCN2 [38] 71.48$\pm$1.70 70.99$\pm$2.45 69.40$\pm$2.10 60.07$\pm$1.73 59.03$\pm$1.79 56.17$\pm$1.16 61.88$\pm$5.17
ASFormer [64] 71.69$\pm$1.91 70.75$\pm$2.17 69.18$\pm$2.75 59.09$\pm$2.64 55.52$\pm$2.29 53.89$\pm$2.29 56.71$\pm$4.60
DiffAct [41] 71.15$\pm$1.32 69.72$\pm$2.16 65.45$\pm$2.22 49.12$\pm$5.08 43.32$\pm$3.52 38.86$\pm$1.78 60.62$\pm$4.21
TSC MiniRocket [12] 61.28$\pm$1.13 60.41$\pm$1.32 57.87$\pm$1.22 62.00$\pm$0.55 61.75$\pm$0.33 58.38$\pm$0.60 62.39$\pm$1.42
TimesNet [60] 67.47$\pm$3.23 65.39$\pm$3.15 63.45$\pm$3.03 59.50$\pm$1.10 57.72$\pm$1.02 55.73$\pm$1.06 50.99$\pm$1.03
PatchTST [47] 51.79$\pm$3.91 55.38$\pm$3.71 52.67$\pm$4.95 58.40$\pm$0.90 56.16$\pm$0.89 53.05$\pm$1.82 58.45$\pm$2.19
NLL SIGUA [24] 67.37$\pm$1.87 65.24$\pm$2.53 63.47$\pm$2.38 54.28$\pm$1.19 53.07$\pm$1.16 51.32$\pm$1.68 53.19$\pm$1.75
UNICON [32] 61.15$\pm$3.72 60.45$\pm$5.22 57.35$\pm$5.57 62.26$\pm$0.66 61.63$\pm$1.22 58.34$\pm$0.99 60.53$\pm$3.96
Sel-CL [39] 63.86$\pm$1.69 62.45$\pm$2.46 61.75$\pm$1.74 63.48$\pm$1.10 63.45$\pm$0.94 61.72$\pm$0.79 60.50$\pm$8.01
TSC & NLL SREA [8] 70.10$\pm$1.28 69.65$\pm$1.22 69.40$\pm$0.85 48.81$\pm$1.09 48.80$\pm$1.16 45.72$\pm$1.56 55.21$\pm$2.25
Scale-T [43] 70.40$\pm$4.14 68.06$\pm$3.03 66.51$\pm$2.98 63.21$\pm$0.85 63.40$\pm$0.73 60.77$\pm$0.99 67.64$\pm$1.42
Con4m 71.28$\pm$2.11 71.27$\pm$2.92 70.04$\pm$2.14 68.02$\pm$1.22 66.31$\pm$1.50 64.31$\pm$1.51 72.00$\pm$1.36
Results of different methods. For fNIRS, TAS models achieve competitive performance compared to Con4m, demonstrating the advantage in modeling contextual data dependency among segments. For Sleep and SEEG data with more ambiguous boundaries, the performance of TAS models deteriorates significantly, and TSC and NLL models slightly outperform TAS models. Benefiting from multi-scale modeling, Scale-T exhibits significantly better performance on the Sleep and SEEG data compared to SREA. Nevertheless, Con4m that fully consider contextual information demonstrate a notable performance improvement (Sleep-0%: $7.15\%$; SEEG: $6.45\%$) in more complex and ambiguous data.
Results of different $r\%$. NLL methods demonstrate close performance degradation as $r\%$ increases from $0\%$ to $20\%$ compared with Con4m. However, with a higher ratio from $20\%$ to $40\%$, SIGUA, UNICON, Sel-CL, SREA, and Scale-T show averaged $3.01\%$, $5.23\%$, $1.92\%$, $3.34\%$, and $3.22\%$ decrease across fNIRS and Sleep data respectively, while Con4m shows $2.37\%$ degradation. For TSC models, non-deep learning-based MiniRocket shows a more robust performance compared to other TSC models. The performance of PatchTST on fNIRS data exhibits significant instability, possibly due to its tendency to overfit inconsistent labels too quickly. DiffAct in TAS models shows the most sensitive performance to boundary perturbations ($14.44\%$ decrease). The stable performance of Con4m indicates that our proposed training framework can effectively harmonize inconsistent labels.
Results of symmetric disturbance. We also corrupt the labels with symmetric disturbance based on segmented labels $y$ rather than raw MVD labels, which is commonly employed in the NLL works [58, 39, 28] of the image classification domain. As shown in Figure 4(b), compared to our novel boundary disturbance, Con4m exhibits stronger robustness to symmetric disturbance. Even with the $20\%$ disturbance ratio, Con4m treats it as a form of data augmentation, resulting in improved performance. This indicates that overcoming more challenging boundary disturbance aligns better with the nature of time series data.
Figure 4: Comparison results of label substitution and symmetric disturbance experiments.
4.3 Label Substitution Experiment
Since ambiguous boundaries are inherent to SEEG data and the majority voting procedure is costly, we limit this procedure to only one high-quality testing group in the label disturbance experiment. Besides, on the SEEG data, Con4m modifies approximately $10\%$ of the training labels, which is a significant proportion. Therefore, it is necessary to further evaluate the effectiveness of our label harmonization process on SEEG data. Specifically, we train the TSC baselines based on the harmonized labels generated by Con4m and observe to what extent the performance of TSC models is improved. As shown in Figure 4(a), PatchTST and TimesNet, employing deep learning architectures, are more susceptible to label inconsistency, so they obtain more significant performance improvement ($4.11\%$ and $7.53\%$). Unlike modified PatchTST that considers the contextual data information across consecutive segments, TimesNet only focuses on the independent segments, thus having a more dramatic improvement. In contrast, MiniRocket achieves only a $1.68\%$ increase, indicating that MiniRocket is more robust with a non-deep learning-based simple random feature mapping.
Table 3: Comparison with model ablations in the $F_{1}$ score (%) in inconsistent scenarios. The best results are in bold and we underline the second best results. The worst results are denoted in italics.
Dataset Model Preserve one Remove one
+ Con-T + Coh-P + Cur-L - Con-T - Coh-P - Cur-L - Fit - $\eta$ Con4m
$r\%$ Acc. $F_{1}$ Acc. $F_{1}$ Acc. $F_{1}$ Acc. $F_{1}$ Acc. $F_{1}$ Acc. $F_{1}$ Acc. $F_{1}$ Acc. $F_{1}$ Acc. $F_{1}$
Sleep 20 65.97 65.05 65.76 65.10 65.31 64.76 65.73 65.53 65.84 65.07 65.85 65.43 66.06 65.28 62.02 59.97 66.61 66.31
40 63.94 62.67 64.42 62.76 63.69 62.23 64.44 63.05 64.23 63.03 64.89 63.07 64.69 63.22 61.93 57.98 65.34 64.31
SEEG - 71.68 67.85 71.69 69.04 71.32 67.22 73.85 70.59 72.41 68.26 74.17 71.18 73.47 70.63 70.70 66.04 74.60 72.00
4.4 Ablation Experiment
We introduce two types of model variations. (1) Preserve only one module. We preserve only the Con-Transformer (Con-T), Coherent Prediction (Coh-P), or Curriculum Learning (Cur-L) module separately. (2) Remove only one component. In addition to removing the above three modules, we also remove the function fitting component (-Fit) and $\eta$ ($E_{\eta}=0$) to verify the necessity of prediction behavior constraint and progressively updating labels.
As shown in Table 3, when keeping one module, +Coh-P achieves the best performance with an averaged $2.78\%$ decrease in $F_{1}$ score, indicating that introducing the contextual label information are most effective for MVD. The utility of each module varies across datasets. For example, for Sleep data, the Con-T contributes more to performance improvement compared to the Cur-L module, while the opposite phenomenon is observed for SEEG data. As for removing one component, even when we only remove the Tanh function fitting, the $F_{1}$ score significantly decreases $1.72\%$ on average. On the Sleep-$20\%$ and SEEG data, the drop caused by -Fit is more significant than that caused by some other modules. Moreover, the model variation -$\eta$ achieves the worst results ($9.23\%$ decrease in $F_{1}$). The results imply that during early training stages, the model tends to learn the consistent parts of the raw labels. Premature use of unreliable predicted labels as subsequent training supervision signals leads to model poisoning and error accumulation.
Figure 5: Case study for a continuous time interval in SEEG testing set.
4.5 Case Study
We present a case study to provide a specific example that illustrates how Con4m works for MVD in Figure 5. We show comparative visualization results for the predictions in a continuous time interval in the SEEG testing set. In SEEG data, we assign the label of normal segments as 00 and that of seizures as $1$. As the figure shows, Con4m demonstrates a more coherent narrative by constraining the prediction behavior and aligning with the contextual data information. In contrast, Scale-T, Sel-CL and MiniRocket exhibit noticeably interrupted and inconsistent predictions. MS-TCN2 fails to identify normal segments. More impressively, Con4m accurately identifies the consistent boundary within the time interval spanning across two classes. This verifies that the label consistency framework can harmonize the boundaries more effectively. Refer to Appendix H for more cases.
5 Conclusion and Discussion
In this work, we focus on the raw time series MVD for segmented time series classification (TSC) tasks, demonstrating unique challenges that are overlooked by existing mainstream TSC models. We first formally demonstrate that valuable contextual information enhances the discriminative power of classification instances. Based on the insights, we introduce contextual prior knowledge of data locality and label coherence to guide the model to focus on contextual information more conducive to discriminating consecutive segments in segmented TSC tasks. Leveraging effective contextual information, a label consistency learning framework Con4m is proposed to progressively harmonize inconsistent labels during training. Extensive experiments validate the superior performance achieved by Con4m and highlight the effectiveness of the proposed consistent label training framework. Our work still has some limitations. We have solely focused on analyzing and designing end-to-end supervised models. Further exploration of large-scale models would be challenging yet intriguing. When faced with more diverse label behaviors, the function fitting module needs to engage in more selection and design of basis functions. Nevertheless, our work brings new insights to the TSC domain, re-emphasizing the importance of the inherent temporal dependence of time series.
Appendix A Details of Related Works
Time series classification (TSC). TSC has become a popular field in various applications with the exponential growth of available time series data in recent years. In response, researchers have proposed numerous algorithms [30]. High accuracy in TSC is achieved by classical algorithms such as Rocket and its variants [11, 12], which use random convolution kernels with relatively low computational cost, as well as ensemble methods like HIVE-COTE [40], which assign weights to individual classifiers.
Moreover, the flourishing non-linear modeling capacity of deep models has led to an increasing prevalence of TSC algorithms based on deep learning. Various techniques are utilized in TSC: RNN-based methods [50, 13] capture temporal changes through state transitions; MLP-based methods [19, 61] encode temporal dependencies into parameters of the MLP layer; and the latest method TimesNet [60] converts one-dimensional time series into a two-dimensional space, achieving state-of-the-art performance on five mainstream tasks. Furthermore, Transformer-based models [63, 9] with attention mechanism have been widely used.
The foundation of our work lies in these researches, including the selection of the backbone and experimental setup. However, mainstream TSC models [46, 18] are often designed for publicly available datasets [3, 10] based on the i.i.d. samples, disregarding the inherent contextual dependencies between classified segments in MVD. Although some time series models [51, 47] use patch-by-patch technique to include contextual information, they are partially context-aware since they only model the data dependencies within each time segment, ignoring the dependencies between consecutive segments.
Noisy label learning (NLL). NLL is an important and challenging research topic in machine learning, as real-world data often rely on manual annotations prone to errors. Early works focus on statistical learning [1, 37, 5]. Researches including Sukhbaatar et al. [53] launch the era of noise-labeled representation learning.
The label noise transition matrix, which represents the transition probability from clean labels to noisy labels [25], is an essential tool. Common techniques for loss correction include forward and backward correction [49], while masking invalid class transitions with prior knowledge is also an important method [22]. Adding an explicit or implicit regularization term in objective functions can reduce the model’s sensitivity to noise, whereas re-weighting mislabeled data can reduce its impact on the objective [2, 65, 42]. Other methods involve training on small-loss instances and utilizing memorization effects. MentorNet [31] pretrains a secondary network to choose clean instances for primary network training. Co-teaching [23] and Co-teaching+ [66], as sample selection methods, introduce two neural networks with differing learning capabilities to train simultaneously, which filter noise labels mutually. The utilization of contrastive learning has emerged as a promising approach for enhancing the robustness in the context of classification tasks of label correction methods [39, 68, 28].
These works primarily focus on handling noisy labels. And ensuring overall label consistency by modifying certain labels is crucial for MVD. To the best of our knowledge, Scale-teaching [43] and SREA [8] are the only NLL works specifically designed for time series. Scale-teaching designs a fine-to-coarse cross-scale fusion mechanism for learning discriminative patterns by utilizing time series at different scales to train multiple DNNs simultaneously. SREA trains a classifier and an autoencoder with a shared embedding representation, progressively self-relabeling mislabeled data samples in a self-supervised manner. However, they still face the issue of overlooking contextual dependencies across consecutive time segments.
Curriculum learning (CL). Bengio et al. [6] propose CL, which imitates human learning by starting with simple samples and progressing to complicated ones. Based on this notion, CL can denoise noisy data since learners are encouraged to train on easier data and spend less time on noisy samples [20, 56]. Current mainstream approaches include Self-paced Learning [36], where students schedule their learning, Transfer Teacher [59], based on a predefined training scheduler; and RL Teacher [21, 44], which incorporates student feedback into the framework. The utilization of CL proves to be particularly advantageous in situations involving changes in the training labels. Hence, this technique is utilized to enhance the harmonization process of boundary labels from MVD in a more stable manner.
Temporal action segmentation (TAS). TAS is a critical task in video understanding and analysis. It involves segmenting an untrimmed video sequence into meaningful temporal segments and assigning a predefined action label to each segment [16]. The majority works on TAS typically take visual feature vectors, either hand-crafted (IDT) [55] or extracted from an off-the-shelf CNN backbone (I3D) [7], as input for each frame. TAS uses sequential modeling to incorporate temporal dependencies and sequential context for improved TAS accuracy.
While TAS shares some similarities with segmented TSC, there still exist some problems. Time series data often have high sampling rates, however, it is impossible to input excessively long sequences into TAS models designed for video data. Moreover, unlike the I3D features [7] used as input in these works, time series lack a unified pretrained model for feature extraction. Besides, the changes between adjacent video frames are smooth and continuous, whereas time series are more variable and exhibit more ambiguous boundaries. Also, TAS works focus on modeling the dependency of instances from a data perspective without explicitly leveraging contextual label information.
Appendix B Implementation Details of Prediction Behavior Constraint
To fit the hyperbolic tangent function (Tanh), we use the mean squared error (MSE) loss function. In practice, we use the Adam optimizer with a learning rate of $0.1$ to optimize the trainable parameters. The maximum number of iterations is set to $100$, and the tolerance value for stopping the fitting process based on loss change is set to $61e-$ 6. Sequences belonging to one minibatch are parallelized to fit their respective Tanh functions. To adapt to the value range of the standard Tanh function, we rescale the sequential predictions to $[-1,1]$ before fitting.
However, it can be difficult to achieve a good fit when fitting with the Tanh function. Specifically, random initialization may fail to fit the sequential values properly when a long time series undergoes a state transition near the boundary. For example, as Figure 6(a) shows, we fit a sequence in which only the last value is $1$. We set all default initial parameters as $1$ and fit it. It can be observed that the fitting function cannot properly fit the trend and will mislabel the last point.
Figure 6: Cases for Tanh fitting.
Appropriate parameter initialization is needed to avoid excessive bias. After careful observation, we find that parameter $k$ controls the slope at the transition part of Tanh, and parameter $b$ controls the abscissa at the transition point. In the process, all fitting values are assigned with uniform abscissa values. Therefore, we calculate the maximum difference between adjacent values and the corresponding position in the entire sequence. And these two values are assigned to parameters $k$ and $b$, respectively. This allows us to obtain suitable initial parameters and avoid getting trapped in local optima or saddle points during function fitting. Formally, given the $L$-length input sequence $\tilde{p}$, we initialize parameters $k$ and $b$ as follows:
$di$ $=\left[\tilde{p}_{i+1}-\tilde{p}_{i}\right]_{i\in\{1,\dots,L-1\}},$
$k,b$ $=\max{\left(\text{Abs}(di)\right)},\arg\max{\left(\text{Abs}(di)\right)},$
$k$ $=k\times\text{Sign}(di[b]),$
$b$ $=-\left(b-\lfloor L/2\rfloor+0.5\right),$
where $\text{Abs}(\cdot)$ and $\text{Sign}(\cdot)$ denote the absolute value function and sign function respectively. $di$ is the difference vector. After proper initialization, as Figure 6(b) shows, we can obtain more accurate fitting results to reduce the probability of mislabeling. We also show some other cases (Figure 6(c)(d)) for the fitting results to verify the effectiveness of the fitting process we propose.
Appendix C Hyperparameter Analysis
Figure 7: Visualization of data division in curriculum learning and hyperparameter analysis of $E_{\eta}$.
The dynamic weighting factor $\eta$ is introduced to progressively update the labels, preventing the model from overly relying on its own predicted labels too early. To validate the utility of $\eta$ and determine an appropriate linear growth epoch $E_{\eta}$, we conduct the hyperparameter search experiment on SEEG data. As shown in Figure 7(b), with smaller $E_{\eta}$ (corresponding to a higher growth rate), there is a significant improvement in model performance. This aligns with our motivation that during the early stage of model training, the primary objective is to better fit the original labels. At this stage, the model’s own predictions are unreliable. If the predicted results are used as training labels too early in subsequent epochs, the model would be adversely affected by its own unreliability. On the other hand, excessively large $E_{\eta}$ leads to a slower rate of label updates, making it more challenging for the model to timely harmonize inconsistent labels. Nonetheless, considering the impact of variance, the model exhibits robustness to slightly larger $E_{\eta}$. In this work, we uniformly use $E_{\eta}=30$ as the default value.
Appendix D Details of Datasets
fNIRS. All signals are sampled at a frequency of $5.2$Hz. At each time step, they record $8$ real-valued measurements, with each measurement corresponding to $2$ concentration changes (oxyhemoglobin and deoxyhemoglobin), $2$ types of optical data (intensity and phase), and $2$ spatial positions on the forehead. Each measurement unit is a micromolar concentration change per liter of tissue (for oxy-/deoxyhemoglobin). They label each part of the active experiment with one of four possible levels of $n$-back working memory intensity (00-back, $1$-back, $2$-back, or $3$-back). More specifically, in an $n$-back task, the subject receives $40$ numbers in sequence. If a number matches the number $n$ steps back, the subject is required to respond accordingly. There are $16$ rounds of tasks, with a $20$-second break between each task. Following Huang et al. [27], we only apply classification tasks for 00-back and $2$-back tasks in our work. Therefore, we only extract sequences for 00-back and $2$-back tasks and concatenate them in chronological order.
Sleep. The Sleep-EDF database records PolySomnoGraphic sleep data from $197$ subjects, including EEG, EOG, chin EMG, and event markers. Some data also includes respiration and temperature-related signals. The database contains two studies: the Sleep Cassette study and the Sleep Telemetry study. The former records approximately $40$ hours of sleep from two consecutive nights, while the latter records around $18$ hours of sleep. Well-trained technicians manually score the corresponding sleep graphs according to the Rechtschaffen and Kales manual. The data is labeled in intervals of $30$ seconds, with each interval being marked as one of the eight possible stages: W, R, 1, 2, 3, 4, M, or ?. In our work, we utilize only the data from the Sleep Cassette study, and retain only the signals from the EEG Fpz-Cz channel and EOG horizontal channel. The EEG and EOG signals were sampled at a frequency of $100$Hz. Following Kemp et al. [33], we remove the labels for stages ? and M from the data, and merge stages 3 and 4, resulting in a $5$-classification task.
SEEG. The private SEEG data records brain signals indicative of suspected pathological tissue within the brains of seizure patients. They are anonymously collected from a top hospital we cooperate with. For a patient suffering from epilepsy, $4$ to $11$ invasive electrodes with $52$ to $153$ channels are used for recording signals. In total, we have collected $847$ hours of SEEG signals with a high frequency ($1,000$Hz or $2,000$Hz) and a total capacity of $1.2$TB. Professional neurosurgeons help us label the seizure segments for each channel. Before sampling for the database, we remove the bad channels marked by neurosurgeons. Then we uniformly downsample the data to $250$Hz and use a low-pass filter to process the data with a cutoff frequency of $30$Hz. Finally, we normalize and sample the intervals for each channel respectively.
Appendix E Implementation Details of Baselines
  • SREA [8]: This time series classification model with noisy labels jointly trains a classifier and an autoencoder with shared embedding representations. It gradually corrects the mislabelled data samples during training in a self-supervised fashion. We use the default model architecture from the source code provided by the author (https://github.com/Castel44/SREA).
  • Scale-teaching [43]: This work designs a fine-to-coarse cross-scale fusion mechanism for learning discriminative patterns by utilizing time series at different scales to train multiple DNNs simultaneously. It uses well-learned multi-scale time series embeddings for noise label correction at sample feature levels. We modify the code to match our datasets based on the code provided by the author (https://github.com/qianlima-lab/Scale-teaching).
  • SIGUA [24]: This model adopts gradient descent on good data as usual, and learning-rate-reduced gradient ascent on bad data, thereby trying to reduce the effect of noisy labels. We modify the network for time series data based on the open source code provided by SREA, using the code from the author (https://github.com/bhanML/SIGUA).
  • UNICON [32]: UNICON introduces a Jensen-Shannon divergence-based uniform selection mechanism and uses contrastive learning to further combat the memorization of noisy labels. We modify the model for time series data according to the code provided by the author (https://github.com/nazmul-karim170/UNICON-Noisy-Label)
  • Sel-CL [39]: Selective-Supervised Contrastive Learning (Sel-CL) is a latest baseline model in the field of computer vision. It selects confident pairs out of noisy ones for supervised contrastive learning (Sup-CL) without knowing noise rates. We modify the code for time series data, based on the source code provided by the author (https://github.com/ShikunLi/Sel-CL)
  • MiniRocket [12]: Rocket [11] achieves state-of-the-art accuracy for time series classification by transforming input time series using random convolutional kernels, and using the transformed features to train a linear classifier. MiniRocket is a variant of Rocket that improves processing time, while offering essentially the same accuracy. We use the code interface from the sktime package (https://github.com/sktime/sktime).
  • TimesNet [60]: This model focuses on temporal variation modeling. With TimesBlock, it can discover the multi-periodicity adaptively and extract the complex temporal variations from transformed 2D tensors by a parameter-efficient inception block. We use the code from the TSlib package (https://github.com/thuml/Time-Series-Library).
  • PatchTST [47]: This is a self-supervised representation learning framework for multivariate time series by segmenting time series into subseries level patches, which are served as input tokens to Transformer with channel-independence. We modify the code to achieve classification for each patch, based on the source code from the Time Series Library (TSlib) package (https://github.com/thuml/Time-Series-Library).
  • MS-TCN2 [38]: This work proposes two multi-stage architectures for the temporal action segmentation task. While the first stage generates an initial prediction, this prediction is iteratively refined by the higher stages. Instead of the commonly used temporal pooling, they use dilated convolutions to increase the temporal receptive field. We modify the code for time series data, based on the source code provided by the author (https://github.com/sj-li/MS-TCN2).
  • ASFormer [64]: ASFormer is a Transformer-based model for action segmentation tasks. It explicitly brings in the inductive priors of local connectivity and applies a pre-defined hierarchical representation pattern to handle long input sequences. The decoder is also carefully designed to refine the initial prediction from the encoder. We modify the code for time series data, based on the source code provided by the author (https://github.com/ChinaYi/ASFormer).
  • DiffAct [41]: In this work, action predictions are iteratively generated from random noise with input video features as conditions. It also devises a unified masking strategy for the conditioning inputs to enhance the modeling of striking characteristics of human actions. We modify the code for time series data, based on the source code provided by the author (https://github.com/Finspire13/DiffAct).
Appendix F Implementation Details of Con4m
The non-linear encoder $g_{\text{enc}}$ used in Con4m is composed of three 1-D convolution layers. The number of kernels vary across different data and you can find corresponding parameters in the default config file of our source code. We construct the Con-Transformer based on the public codes implemented by HuggingFace [^1] https://github.com/huggingface/transformers/blob/v4.25.1/src/transformers/models/bert/modeling_bert.py. We set $d$=$128$ and the dimension of intermediate representations in FFN module as $256$ for all experiments. The number of heads and dropout rate are set as $8$ and $0.1$ respectively. Since we observe that one-layer Con-Attention can fit the data well, we do not stack more layers to avoid overfitting. Note that Con4m consists of two Con-Transformers and we use two Con-Attention layers. The model is optimized using Adam optimizer [34] with a learning rate of $1e$-$3$ and weight decay of $1e$-$4$, and the batch size is set as $64$. We build our model using PyTorch 2.0.0 [48] with CUDA 11.8. And the model is trained on a workstation (Ubuntu system 20.04.5) with 2 CPUs (AMD EPYC 7H12 64-Core Processor) and 8 GPUs (NVIDIA GeForce RTX 3090). You can find more technical details in our source code attached in the supplementary materials.
TAS TSC NLL TSC&NLL
$r\%$ Exp MS- TCN2 AS Former DiffAct Mini Rocket Times Net Patch TST SI- GUA UNI- CON Sel- CL SREA Scale- T Con4m
0 1 68.61 69.76 71.11 61.37 60.73 51.07 64.75 63.85 63.95 69.13 66.07 68.55
2 70.50 71.85 68.62 62.45 68.25 48.21 67.55 57.17 65.49 70.29 74.26 71.64
3 70.21 70.08 71.42 60.96 66.38 54.23 65.56 61.71 61.44 67.14 68.99 70.51
4 71.89 73.38 72.25 62.24 69.73 55.07 68.83 60.74 63.23 70.23 71.64 72.65
5 71.87 72.91 72.80 61.35 66.46 57.11 67.96 54.87 63.21 70.13 75.85 68.55
6 72.53 73.12 71.29 61.79 69.58 57.22 70.12 58.96 64.66 69.66 70.50 72.99
7 69.08 68.08 70.77 60.11 62.64 50.46 64.03 62.54 60.13 70.55 62.42 70.63
8 72.29 72.44 72.13 62.90 69.57 49.75 69.15 63.05 65.31 70.90 66.81 73.36
9 72.52 72.22 69.47 58.78 67.23 48.86 68.41 66.77 63.45 70.29 71.88 72.60
10 74.48 70.20 73.02 60.75 71.17 55.63 68.24 59.40 65.47 71.59 70.60 69.38
11 70.71 71.23 70.09 60.71 66.64 48.51 65.95 57.84 65.24 69.17 76.77 69.42
12 73.08 75.02 70.81 61.93 71.30 45.40 67.89 66.88 64.70 72.17 68.95 75.14
Avg 71.48 71.69 71.15 61.28 67.47 51.79 67.37 61.15 63.86 70.10 70.40 71.28
Std 1.70 1.91 1.32 1.13 3.23 3.91 1.87 3.72 1.69 1.28 4.14 2.11
20 1 67.10 70.19 72.07 59.22 62.74 49.81 61.42 64.38 60.91 68.32 65.55 69.48
2 71.26 70.42 69.61 60.10 67.31 58.02 63.38 63.30 64.00 70.40 68.36 72.94
3 72.18 66.46 71.39 59.52 60.19 54.99 64.27 51.67 57.25 70.16 64.44 72.06
4 71.19 71.74 71.53 63.15 66.04 62.28 67.91 62.23 61.45 69.78 66.18 73.56
5 71.85 72.38 69.18 62.04 67.66 55.29 66.07 56.26 64.22 68.45 66.59 71.41
6 72.03 69.99 69.93 61.58 68.32 57.22 67.09 62.59 65.45 70.34 68.71 72.08
7 69.94 66.99 68.32 59.15 59.02 53.13 60.97 49.65 59.18 67.69 71.55 62.68
8 69.19 72.09 64.24 60.33 66.27 56.57 63.72 66.42 63.28 70.85 75.03 71.59
9 74.44 72.21 69.70 59.41 66.73 52.13 67.11 62.28 61.54 68.15 66.94 71.84
10 72.15 72.79 69.91 60.77 69.07 56.96 68.27 59.87 65.18 71.29 64.83 71.72
11 66.50 73.26 68.58 58.87 65.17 49.62 64.83 63.95 63.32 69.48 68.88 72.08
12 74.10 70.41 72.16 60.79 66.22 58.49 67.87 62.81 63.62 70.93 69.65 73.76
Avg 70.99 70.75 69.72 60.41 65.39 55.38 65.24 60.45 62.45 69.65 68.06 71.27
Std 2.45 2.17 2.16 1.32 3.15 3.71 2.53 5.22 2.46 1.22 3.03 2.92
40 1 68.74 65.53 65.47 57.21 62.93 51.39 60.63 52.63 61.98 69.37 62.03 65.90
2 67.96 70.88 64.39 58.85 62.10 50.27 59.84 45.74 62.50 69.43 68.10 71.91
3 68.25 68.40 68.79 58.30 60.06 44.09 65.00 54.70 59.58 69.12 64.79 71.05
4 70.76 71.54 66.20 59.23 68.56 58.48 67.18 63.46 64.25 68.84 66.90 70.68
5 72.62 66.59 64.44 57.05 59.96 54.44 64.60 63.00 58.31 68.49 67.67 71.55
6 71.36 72.64 61.00 58.43 66.76 53.18 64.20 59.95 61.72 69.85 68.12 72.75
7 67.31 67.94 68.49 56.25 60.06 49.93 61.58 52.33 60.33 69.19 59.47 66.69
8 66.67 66.67 65.84 58.26 68.32 53.20 67.27 60.76 63.91 70.49 68.52 68.50
9 66.59 67.74 66.30 56.95 63.86 61.90 61.80 65.20 62.82 68.42 67.82 69.88
10 69.56 66.67 65.55 55.78 62.01 49.37 63.19 55.84 61.04 70.16 66.68 69.00
11 71.66 72.80 62.43 58.34 62.78 57.77 62.12 57.92 61.68 68.36 68.64 70.59
12 71.32 72.74 66.53 59.81 63.96 48.00 64.21 56.73 62.93 71.05 69.38 72.02
Avg 69.40 69.18 65.45 57.87 63.45 52.67 63.47 57.35 61.75 69.40 66.51 70.04
Std 2.10 2.75 2.22 1.22 3.03 4.95 2.38 5.57 1.74 0.85 2.98 2.14
Table 4: Full results of the label disturbance experiment on fNIRS data. The best results are in bold and we underline the second best results. The worst results are denoted in italics.
TAS TSC NLL TSC&NLL
$r\%$ Exp MS- TCN2 AS Former DiffAct Mini Rocket Times Net Patch TST SI- GUA UNI- CON Sel- CL SREA Scale- T Con4m
0 1 59.46 55.89 51.79 62.16 58.73 58.42 54.79 62.41 63.49 48.95 63.05 68.80
2 59.47 58.37 39.55 61.14 59.72 58.60 52.69 62.49 62.87 46.93 62.70 67.63
3 60.72 59.60 52.65 62.74 60.76 59.44 56.19 62.62 65.47 49.38 64.91 69.29
4 61.41 58.65 47.42 61.64 58.23 59.22 53.51 61.01 62.88 48.55 62.94 66.66
5 57.26 58.17 50.62 62.20 60.80 57.45 54.36 62.89 63.82 48.82 63.07 66.61
6 62.12 63.86 52.68 62.10 58.76 57.25 54.12 62.11 62.37 50.24 62.61 69.11
Avg 60.07 59.09 49.12 62.00 59.50 58.40 54.28 62.26 63.48 48.81 63.21 68.02
Std 1.73 2.64 5.08 0.55 1.10 0.90 1.19 0.66 1.10 1.09 0.85 1.22
20 1 59.81 57.52 37.81 61.86 58.07 56.82 53.73 62.75 64.41 49.80 62.29 67.07
2 58.06 51.43 48.61 61.51 57.44 55.50 51.04 62.68 63.58 47.56 63.78 64.25
3 61.31 55.38 42.49 62.35 56.10 57.03 54.51 61.44 64.58 49.30 64.28 68.50
4 56.09 57.82 44.68 61.61 57.23 56.77 53.12 59.39 62.33 47.65 63.43 65.25
5 59.07 55.79 42.49 61.75 58.99 54.78 52.83 61.92 63.28 48.18 63.82 65.90
6 59.85 55.21 43.84 61.43 58.47 56.05 53.18 61.60 62.51 50.28 62.81 66.86
Avg 59.03 55.52 43.32 61.75 57.72 56.16 53.07 61.63 63.45 48.80 63.40 66.31
Std 1.79 2.29 3.52 0.33 1.02 0.89 1.16 1.22 0.94 1.16 0.73 1.50
40 1 56.51 56.67 37.69 58.62 57.20 55.98 52.10 58.17 61.54 47.23 61.15 65.38
2 54.02 54.82 38.80 57.96 55.26 52.60 50.08 58.12 61.64 44.56 60.65 64.27
3 56.53 51.22 36.93 59.18 55.30 52.12 53.85 59.63 63.27 47.98 62.10 65.36
4 55.76 56.15 39.68 58.68 54.36 54.31 52.21 57.58 61.59 45.53 61.09 65.69
5 57.14 52.19 41.95 57.47 56.80 50.80 49.48 57.16 61.28 44.03 60.53 61.82
6 57.06 52.31 38.13 58.36 55.44 52.50 50.18 59.40 61.01 45.00 59.08 63.33
Avg 56.17 53.89 38.86 58.38 55.73 53.05 51.32 58.34 61.72 45.72 60.77 64.31
Std 1.16 2.29 1.78 0.60 1.06 1.82 1.68 0.99 0.79 1.56 0.99 1.51
Table 5: Full results of the label disturbance experiment on Sleep data. The best results are in bold and we underline the second best results. The worst results are denoted in italics.
TAS TSC NLL TSC&NLL
$r\%$ Exp MS- TCN2 AS Former DiffAct Mini Rocket Times Net Patch TST SI- GUA UNI- CON Sel- CL SREA Scale- T Con4m
- 1 65.86 51.41 64.82 62.11 50.25 58.51 52.09 60.64 62.57 53.09 66.80 72.26
2 63.74 59.68 60.63 61.12 50.55 60.60 52.27 64.43 51.65 57.57 66.85 73.21
3 56.03 59.04 56.41 63.92 52.17 56.22 55.21 56.51 67.27 54.99 69.28 70.52
Avg 61.88 56.71 60.62 62.39 50.99 58.45 53.19 60.53 60.50 55.21 67.64 72.00
Std 5.17 4.60 4.21 1.42 1.03 2.19 1.75 3.96 8.01 2.25 1.42 1.36
Table 6: Full results of the label disturbance experiment on SEEG data. The best results are in bold and we underline the second best results. The worst results are denoted in italics.
Appendix G Full Results
The full results of the label disturbance experiment are listed in Table 45 and 6. For fNIRS, we first divide the data into $4$ groups by subjects and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct cross-validation experiments. Therefore, there are $C_{4}^{2}\times C_{2}^{1}=12$ experiments in total. Similarly, we divide the Sleep data into $3$ groups and follow the $1$-$1$-$1$ experimental setting. Therefore, we carry out $C_{3}^{1}\times C_{2}^{1}=6$ experiments. For SEEG data, we follow the same setting as fNIRS. Notice that for SEEG data, inconsistent labels already exist in the raw data. We obtain a high-quality testing group by using a majority voting procedure to determine the boundaries. Then we leave the testing group aside and only change the validation group to report the mean value of $C_{3}^{2}=3$ experiments. All the experimental results are listed in lexicographical order according to the group name composition. We also report the mean value and standard derivation of all experiments. Specifically, we use the STDEVA to estimate standard deviation based on a sample of data.
Appendix H Case Study
As shown in Figure 8, we present four cases to compare and demonstrate the differences between our proposed Con4m and other baselines. The first two cases involve transitions from a seizure state of label $1$ to a normal state of label 00. The third case consists of entirely normal segments, while the fourth case comprises entirely seizure segments. As illustrated in the figure, Con4m exhibits more coherent narratives by constraining the predictions to align with the contextual information of the data. Moreover, it demonstrates improved accuracy in identifying the boundaries of transition states. In contrast, other baselines exhibit fragmented and erroneous predictions along the time segments. This verifies that Con4m can achieve clearer recognition of boundaries, and it can also make better predictions on the consecutive time segments belonging to the same class.
Figure 8: More cases for continuous time intervals in SEEG testing set.

### v2

@@
 Extensive experiments across multiple datasets validate the effectiveness of Con4m in handling segmented TSC tasks on MVD.
+The source code is available at https://github.com/MrNobodyCali/Con4m.
 footnotetext: $\dagger$ Corresponding author.
 1 Introduction
-Figure 1: (a) In the healthcare domain, different physicians may reach a consensus on the seizure onset, but they have varying annotations regarding the end times of seizure waves.
-(b) Reasonable model predictions exhibit coherence across consecutive segments rather than repeated interruptions.
+Figure 1: (a) Reasonable model predictions exhibit coherence across consecutive segments rather than repeated interruptions.
+(b) In the healthcare domain, different physicians have varying annotations regarding the start and end times of seizure waves.
 (c) Based on the proximity to the boundary, we divide each class sequence into 5 levels, from which an equal number of segments are sampled.
@@
 We take the seizure detection task as an example, in which given the brain signals of epileptic patients, the model should identify whether a segment includes seizure waves or not.
-As illustrated in Figure 1(b), given the sustained nature of seizure onsets, the model predictions for consecutive segments should exhibit coherence, with seizure and normal predictions appearing in continuous and concentrated patterns.
+As illustrated in Figure 1(a), given the sustained nature of seizure onsets, the model predictions for consecutive segments should exhibit coherence, with seizure and normal predictions appearing in continuous and concentrated patterns.
 However, mainstream TSC models [58, 71, 61] focus on the i.i.d.
@@
 In domains with precious labels, different annotators collaboratively contribute annotations.
-The raw annotations of MVD typically include information about the start and end times for each class.
-However, due to inherent ambiguity and a lack of unified quantification standards, differences in annotator experience may lead to inconsistent boundary labels.
-Returning to Figure 1(a), in the seizure detection task, owing to the natural fuzzy transition from seizure onset to completely normal, different physicians have varying experiences regarding when seizure waves terminate.
+The raw annotations of MVD typically include the start and end times for each class.
+However, due to inherent ambiguity and a lack of unified quantification standards, for MVD, the boundaries between states are not clearly defined, or the transitional state itself represent a mixed state.
+Consequently, behind inconsistent labels, there is no artificially defined true label.
+Therefore, our work aims to harmonize this inconsistency as much as possible to reduce the instability of model training and enhance its performance.
+Returning to Figure 1(b), in the seizure detection task, owing to the natural fuzzy transition from seizure onset to completely normal, different physicians have varying experiences regarding when seizure waves terminate.
 Furthermore, inconsistent boundary labeling causes that boundary segments with similar patterns may have opposite labels, leading to unstable model training.
@@
 (1) We are the first to propose a practical consistency learning framework Con4m for the segmented TSC based on the raw MVD.
-(2) By comprehensively integrating prior knowledge from both the data and label perspectives, we guide the model to focus on effective contextual information.
+(2) By comprehensively integrating prior knowledge from the data and label perspectives, we guide the model to focus on effective contextual information.
 Based on context-aware predictions, a progressive harmonization approach for handling inconsistent training labels is designed to yield a more robust model.
-(3) Extensive experiments on two public and one private MVD datasets demonstrate the superior performance of Con4m.
-The Con4m’s ability to harmonize inconsistent labels is further verified by the label substitution experiment and a case study.
+(3) Extensive experiments on three public and one private MVD datasets demonstrate the superior performance of Con4m.
+The Con4m’s ability to harmonize inconsistent labels is further verified by the label substitution experiment and case study.
 2 Theoretical Analysis
@@
 Since the neighbors of boundary segments may belong to different classes, we allow each segment to learn its own scale parameter $\sigma$.
-Formally, as Figure 2 shows, the two-branch Con-Attention in the $l$-th layer is:
+Formally, as Figure 2(a) shows, the two-branch Con-Attention in the $l$-th layer is:
 $Q,K,V_{s},V_{g},\sigma=$ $c^{l-1}W_{Q}^{l},c^{l-1}W_{K}^{l},c^{l-1}W_{V_{s}}^{l},c^{l-1}W_{V_{g}}^{l},c^{l-1}W_{\sigma}^{l},$
@@
 We use the conventional attention mechanism [4] to adaptively fuse $z_{s}^{l}$ and $z_{g}^{l}$.
-Finally, as illustrated in Figure 2, by stacking the multi-head version of Con-Attention layers, we construct Con-Transformer, which serves as the backbone of the continuous encoder of Con4m to obtain final representations $c$.
-Figure 2: Overview of the encoder of Con4m. The leftmost part shows the details of Con-Attention. The right part of the figure shows the architecture of Con-Transformer and the whole encoder of Con4m.
+Finally, as illustrated in Figure 2(a), by stacking the multi-head version of Con-Attention layers, we construct Con-Transformer, which serves as the backbone of the continuous encoder of Con4m to obtain final representations $c$.
+We employ learnable absolute positional encoding (APE) [23] for the input representations.
+Figure 2: Overview of Con4m.
+(a) Overview of continuous contextual representation encoder in Con4m.
+The leftmost part shows the details of Con-Attention.
+The right part of the figure shows the architecture of Con-Transformer and the whole encoder of Con4m.
+(b) Overview of context-aware coherent class prediction and consistent label training framework in Con4m.
+The right part describes the neighbor class consistency discrimination task and the prediction behavior constraint.
+The leftmost part presents the training and inference details for label harmonization.
 3.2 Context-aware Coherent Class Prediction
@@
 The model then aggregates the contextual class predictions based on the discriminator’s outputs, thus making more robust predictions.
-As the left part of Figure 3 shows, we formalize this process as follows:
+As the right part of Figure 2(b) shows, we formalize this process as follows:
 $\hat{R}=\text{SoftMax}\left(\left[\text{MLP}_{2}\left(c_{i}\|c_{j}\right)\right]_{i,j\in\{1,\dots,L\}}\right),\quad\hat{p}=\text{SoftMax}\left(\text{MLP}_{1}\left(c\right)\right),\quad\tilde{p}=\hat{R}_{:,:,1}\hat{p},$
@@
 Therefore, we ensure the monotonicity of predictions across consecutive segments through hard constraints, thereby utilizing contextual label information $\textnormal{y}_{\mathbb{A}_{t}}$ to integrate and refine predictions across these segments.
-As shown in the middle part of Figure 3, for each class in the predictions, there are only four prediction behaviors for consecutive segments, namely high confidence, low confidence, confidence decreasing, and confidence increasing.
+As shown in the middle part of Figure 2(b), for each class in the predictions, there are only four prediction behaviors for consecutive segments, namely high confidence, low confidence, confidence decreasing, and confidence increasing.
 To constrain the behavior, we use function fitting to integrate $\tilde{p}$.
@@
 For the inference stage, we use the average of them as the final coherent predictions, i.e., $\hat{y}=\arg\max{(\hat{p}+\bar{p})/2}$.
-Figure 3: Overview of context-aware coherent class prediction and consistent label training framework in Con4m.
-The left part describes the neighbor class consistency discrimination task and the prediction behavior constraint.
-The rightmost part presents the training and inference details for label harmonization.
 3.3 Label Consistency Training Framework
@@
 The higher the level, the more apparent the inconsistency.
-Therefore, as the right part of Figure 3 shows, during the training stage, Con4m learns the time intervals in order from low to high levels, with a lag gap of $E_{g}=5$ epochs.
+Therefore, as the left part of Figure 2(b) shows, during the training stage, Con4m learns the time intervals in order from low to high levels, with a lag gap of $E_{g}=5$ epochs.
 Harmonizing inconsistent labels.
@@
 The dynamic weighting factor, $\eta$, is used to adjust the degree of label update.
-As Figure 3 shows, $\eta$ linearly increases from 00 to $1$ with $E_{\eta}$ epochs, gradually weakening the influence of the original labels.
+As the left part of Figure 2(b) shows, $\eta$ linearly increases from 00 to $1$ with $E_{\eta}$ epochs, gradually weakening the influence of the original labels.
 Besides, in the initial training stage, the model tends to improve independent predictions.
@@
 Datasets.
-In this work, we use two public [27, 33] and one private MVD data to measure the performance of models.
+In this work, we use three public [31, 7, 37] and one private MVD data to measure the performance of models.
 Specifically, the Tufts fNIRS to Mental Workload [31] data (fNIRS) contains brain activity recordings from adult humans performing controlled cognitive workload tasks.
+The HHAR (Heterogeneity Human Activity Recognition) dataset [7] captures sensor data from multiple smart devices to explore the impact of device heterogeneity on human activity recognition.
 The SleepEDF [37] data (Sleep) contains PolySomnoGraphic sleep records for subjects over a whole night.
@@
 fNIRS 5.2Hz 8 2 68 4 12 4,080 38.46s 4.81s 0.96s 146,880
+HHAR 50Hz 6 6 9 3 6 5,400 60s 4s 2s 156,600
 Sleep 100Hz 2 5 154 3 6 6,000 40s 2.5s 1.25s 186,000
@@
 As shown in Table 1, for fNIRS and SEEG, we divide the subjects into $4$ groups and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct experiments.
-We divide the Sleep dataset into $3$ groups and follow the $1$-$1$-$1$ experimental setting.
-Therefore, we only report the mean values of cross-validation results in the main context.
+We divide the HHAR and Sleep datasets into $3$ groups and follow the $1$-$1$-$1$ experimental setting.
+Notice that SEEG data is derived from real clinical datasets and annotated by multiple experts, resulting in naturally inconsistent labels.
+We employ a voting mechanism which brings annotators together to collectively decide the boundaries to minimize discrepancies in test labels.
+Considering the high cost of this approach, we do not apply it to the training and validation sets.
+Therefore, we leave the test group aside and only change the training and validation groups to conduct cross-validation.
+Finally, we only report the mean values of cross-validation results in the main context.
 See more details and the full results in Appendix G.
@@
 The worst results are denoted in italics.
-Model $r\%$ fNIRS [27] Sleep [33] SEEG
-$0\%$ $20\%$ $40\%$ $0\%$ $20\%$ $40\%$ raw
-TAS MS-TCN2 [38] 71.48$\pm$1.70 70.99$\pm$2.45 69.40$\pm$2.10 60.07$\pm$1.73 59.03$\pm$1.79 56.17$\pm$1.16 61.88$\pm$5.17
-ASFormer [64] 71.69$\pm$1.91 70.75$\pm$2.17 69.18$\pm$2.75 59.09$\pm$2.64 55.52$\pm$2.29 53.89$\pm$2.29 56.71$\pm$4.60
-DiffAct [41] 71.15$\pm$1.32 69.72$\pm$2.16 65.45$\pm$2.22 49.12$\pm$5.08 43.32$\pm$3.52 38.86$\pm$1.78 60.62$\pm$4.21
-TSC MiniRocket [12] 61.28$\pm$1.13 60.41$\pm$1.32 57.87$\pm$1.22 62.00$\pm$0.55 61.75$\pm$0.33 58.38$\pm$0.60 62.39$\pm$1.42
-TimesNet [60] 67.47$\pm$3.23 65.39$\pm$3.15 63.45$\pm$3.03 59.50$\pm$1.10 57.72$\pm$1.02 55.73$\pm$1.06 50.99$\pm$1.03
-PatchTST [47] 51.79$\pm$3.91 55.38$\pm$3.71 52.67$\pm$4.95 58.40$\pm$0.90 56.16$\pm$0.89 53.05$\pm$1.82 58.45$\pm$2.19
-NLL SIGUA [24] 67.37$\pm$1.87 65.24$\pm$2.53 63.47$\pm$2.38 54.28$\pm$1.19 53.07$\pm$1.16 51.32$\pm$1.68 53.19$\pm$1.75
-UNICON [32] 61.15$\pm$3.72 60.45$\pm$5.22 57.35$\pm$5.57 62.26$\pm$0.66 61.63$\pm$1.22 58.34$\pm$0.99 60.53$\pm$3.96
-Sel-CL [39] 63.86$\pm$1.69 62.45$\pm$2.46 61.75$\pm$1.74 63.48$\pm$1.10 63.45$\pm$0.94 61.72$\pm$0.79 60.50$\pm$8.01
-TSC & NLL SREA [8] 70.10$\pm$1.28 69.65$\pm$1.22 69.40$\pm$0.85 48.81$\pm$1.09 48.80$\pm$1.16 45.72$\pm$1.56 55.21$\pm$2.25
-Scale-T [43] 70.40$\pm$4.14 68.06$\pm$3.03 66.51$\pm$2.98 63.21$\pm$0.85 63.40$\pm$0.73 60.77$\pm$0.99 67.64$\pm$1.42
-Con4m 71.28$\pm$2.11 71.27$\pm$2.92 70.04$\pm$2.14 68.02$\pm$1.22 66.31$\pm$1.50 64.31$\pm$1.51 72.00$\pm$1.36
+Model $r\%$ fNIRS [31] HHAR [7] Sleep [37] SEEG
+$0\%$ $20\%$ $40\%$ $0\%$ $20\%$ $40\%$ $0\%$ $20\%$ $40\%$ raw
+TAS MS-TCN2 [42] 71.48 70.99 69.40 69.79 66.72 62.29 60.07 59.03 56.17 61.88
+ASFormer [68] 71.69 70.75 69.18 62.52 60.92 60.77 59.09 55.52 53.89 56.71
+DiffAct [45] 71.15 69.72 65.45 56.76 53.86 50.63 49.12 43.32 38.86 60.62
+TSC MiniRocket [13] 61.28 60.41 57.87 70.34 63.32 59.25 62.00 61.75 58.38 62.39
+TimesNet [64] 67.47 65.39 63.45 72.07 70.19 66.76 59.50 57.72 55.73 50.99
+PatchTST [51] 51.79 55.38 52.67 52.00 45.46 45.69 58.40 56.16 53.05 58.45
+NLL SIGUA [28] 67.37 65.24 63.47 68.94 68.47 67.60 54.28 53.07 51.32 53.19
+UNICON [36] 61.15 60.45 57.35 62.26 61.63 58.34 62.26 61.63 58.34 60.53
+Sel-CL [43] 63.86 62.45 61.75 73.00 72.28 72.81 63.48 63.45 61.72 60.50
+TSC & NLL SREA [9] 70.10 69.65 69.40 68.64 66.02 65.67 48.81 48.80 45.72 55.21
+Scale-T [47] 70.40 68.06 66.51 77.77 76.71 75.97 63.21 63.40 60.77 67.64
+Con4m 71.28 71.27 70.04 80.29 78.59 75.52 68.02 66.31 64.31 72.00
 Results of different methods.
 For fNIRS, TAS models achieve competitive performance compared to Con4m, demonstrating the advantage in modeling contextual data dependency among segments.
-For Sleep and SEEG data with more ambiguous boundaries, the performance of TAS models deteriorates significantly, and TSC and NLL models slightly outperform TAS models.
+For HHAR, Sleep and SEEG data with more ambiguous boundaries, the performance of TAS models deteriorates significantly, and TSC and NLL models slightly outperform TAS models.
 Benefiting from multi-scale modeling, Scale-T exhibits significantly better performance on the Sleep and SEEG data compared to SREA.
-Nevertheless, Con4m that fully consider contextual information demonstrate a notable performance improvement (Sleep-0%: $7.15\%$; SEEG: $6.45\%$) in more complex and ambiguous data.
+Nevertheless, Con4m that fully consider contextual information demonstrate a notable performance improvement (HHAR-0%: 3.24%; Sleep-0%: $7.15\%$; SEEG: $6.45\%$) in more complex and ambiguous data.
 Results of different $r\%$.
 NLL methods demonstrate close performance degradation as $r\%$ increases from $0\%$ to $20\%$ compared with Con4m.
-However, with a higher ratio from $20\%$ to $40\%$, SIGUA, UNICON, Sel-CL, SREA, and Scale-T show averaged $3.01\%$, $5.23\%$, $1.92\%$, $3.34\%$, and $3.22\%$ decrease across fNIRS and Sleep data respectively, while Con4m shows $2.37\%$ degradation.
+However, with a higher ratio from $20\%$ to $40\%$, SIGUA, UNICON, Sel-CL, SREA, and Scale-T show averaged $3.01\%$, $5.23\%$, $1.92\%$, $3.34\%$, and $3.22\%$ decrease across fNIRS and Sleep data, while Con4m shows $2.37\%$ degradation.
 For TSC models, non-deep learning-based MiniRocket shows a more robust performance compared to other TSC models.
 The performance of PatchTST on fNIRS data exhibits significant instability, possibly due to its tendency to overfit inconsistent labels too quickly.
-DiffAct in TAS models shows the most sensitive performance to boundary perturbations ($14.44\%$ decrease).
+DiffAct in TAS models shows the most sensitive performance to boundary perturbations from $0\%$ to $40\%$ across three public data ($13.23\%$ decrease).
 The stable performance of Con4m indicates that our proposed training framework can effectively harmonize inconsistent labels.
@@
 We also corrupt the labels with symmetric disturbance based on segmented labels $y$ rather than raw MVD labels, which is commonly employed in the NLL works [62, 43, 32] of the image classification domain.
-As shown in Figure 4(b), compared to our novel boundary disturbance, Con4m exhibits stronger robustness to symmetric disturbance.
+As shown in Figure 3(a), compared to our novel boundary disturbance, Con4m exhibits stronger robustness to symmetric disturbance.
 Even with the $20\%$ disturbance ratio, Con4m treats it as a form of data augmentation, resulting in improved performance.
 This indicates that overcoming more challenging boundary disturbance aligns better with the nature of time series data.
-Figure 4: Comparison results of label substitution and symmetric disturbance experiments.
+Figure 3: Comparison results of symmetric disturbance and label substitution experiments.
 4.3 Label Substitution Experiment
@@
 Specifically, we train the TSC baselines based on the harmonized labels generated by Con4m and observe to what extent the performance of TSC models is improved.
-As shown in Figure 4(a), PatchTST and TimesNet, employing deep learning architectures, are more susceptible to label inconsistency, so they obtain more significant performance improvement ($4.11\%$ and $7.53\%$).
+As shown in Figure 3(b), PatchTST and TimesNet, employing deep learning architectures, are more susceptible to label inconsistency, so they obtain more significant performance improvement ($4.11\%$ and $7.53\%$).
 Unlike modified PatchTST that considers the contextual data information across consecutive segments, TimesNet only focuses on the independent segments, thus having a more dramatic improvement.
 In contrast, MiniRocket achieves only a $1.68\%$ increase, indicating that MiniRocket is more robust with a non-deep learning-based simple random feature mapping.
+4.4 Ablation Experiment
+We introduce two types of model variations.
+(1) Preserve only one module.
+We preserve only the Con-Transformer (Con-T), Coherent Prediction (Coh-P), or Curriculum Learning (Cur-L) module separately.
+(2) Remove only one component.
+In addition to removing the above three modules, we also remove the function fitting component (-Fit) and $\eta$ ($E_{\eta}=0$) to verify the necessity of prediction behavior constraint and progressively updating labels.
 Table 3: Comparison with model ablations in the $F_{1}$ score (%) in inconsistent scenarios.
@@
 SEEG - 71.68 67.85 71.69 69.04 71.32 67.22 73.85 70.59 72.41 68.26 74.17 71.18 73.47 70.63 70.70 66.04 74.60 72.00
-4.4 Ablation Experiment
-We introduce two types of model variations.
-(1) Preserve only one module.
-We preserve only the Con-Transformer (Con-T), Coherent Prediction (Coh-P), or Curriculum Learning (Cur-L) module separately.
-(2) Remove only one component.
-In addition to removing the above three modules, we also remove the function fitting component (-Fit) and $\eta$ ($E_{\eta}=0$) to verify the necessity of prediction behavior constraint and progressively updating labels.
 As shown in Table 3, when keeping one module, +Coh-P achieves the best performance with an averaged $2.78\%$ decrease in $F_{1}$ score, indicating that introducing the contextual label information are most effective for MVD.
@@
 Premature use of unreliable predicted labels as subsequent training supervision signals leads to model poisoning and error accumulation.
+4.5 Case Study
 Figure 4: Case study for a continuous time interval in SEEG testing set.
-4.5 Case Study
+The C-score, introduced by the ClaSP model [19], assessing the ability of models to recognize segmentation boundaries by measuring the trade-off between precision (correctly identified change points) and recall (finding all true change points).
 We present a case study to provide a specific example that illustrates how Con4m works for MVD in Figure 4.
@@
 More impressively, Con4m accurately identifies the consistent boundary within the time interval spanning across two classes.
+We also utilize the C-score proposed by the ClaSP model [19] to assess the segmentation capability of models.
+The C-score reflects the ability of models to recognize segmentation boundaries by measuring the trade-off between precision (correctly identified change points) and recall (finding all true change points), ensuring the model captures meaningful transitions without over-segmenting or missing important splits.
+We compute the scores for each model across three sets of experiments on SEEG data and take the average.
+Notably, Con4m outperforms the other models significantly, while MS-TCN2, specifically designed for segmentation tasks, also achieves impressive scores.
 This verifies that the label consistency framework can harmonize the boundaries more effectively.
@@
 Further exploration of large-scale models would be challenging yet intriguing.
+Con4m is a combination of segmentation and classification, both of which are fully supervised.
+Exploring its application in unsupervised segmentation tasks is worthwhile.
 When faced with more diverse label behaviors, the function fitting module needs to engage in more selection and design of basis functions.
 Nevertheless, our work brings new insights to the TSC domain, re-emphasizing the importance of the inherent temporal dependence of time series.
+Acknowledgments and Disclosure of Funding
+This work was partially supported by National Natural Science Foundation of China (No. 62322606, No. 62441605).
 Appendix A Details of Related Works
@@
 Therefore, we only extract sequences for 00-back and $2$-back tasks and concatenate them in chronological order.
+HHAR.
+The HHAR (Heterogeneity Human Activity Recognition) dataset [7] captures sensor data from multiple smart devices to explore the impact of device heterogeneity on human activity recognition.
+It involves measurements from accelerometers and gyroscopes, which are two common motion sensors found in smartphones and smartwatches.
+The dataset includes six types of human activities: ‘Biking’, ‘Sitting’, ‘Standing’, ‘Walking’, ‘Stairs Up’, and ‘Stairs Down’.
+For each activity, both 3-axis accelerometer and 3-axis gyroscope readings are recorded, resulting in six key features per time step.
+The data was collected from five different device types, consisting of four smartwatches and eight smartphones, across nine participants.
+Each device samples data at its highest possible frequency, generally around $50$Hz.
+For data preprocessing, the accelerometer and gyroscope signals are aligned based on their timestamps, normalized within each channel, and divided by the devices, following the procedure outlined in Gagnon-Audet et al.
+[21].
+To ensure balanced samples, we combined data from the Galaxy S3 Mini, LG watch, and Gear watch into a single group.
 Sleep.
@@
 And the model is trained on a workstation (Ubuntu system 20.04.5) with 2 CPUs (AMD EPYC 7H12 64-Core Processor) and 8 GPUs (NVIDIA GeForce RTX 3090).
-You can find more technical details in our source code attached in the supplementary materials.
+Appendix G Full Results
+The full results of the label disturbance experiment are listed in Table 456 and 7.
+For fNIRS, we first divide the data into $4$ groups by subjects and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct cross-validation experiments.
+Therefore, there are $C_{4}^{2}\times C_{2}^{1}=12$ experiments in total.
+Similarly, we divide the HHAR and Sleep data into $3$ groups and follow the $1$-$1$-$1$ experimental setting.
+Therefore, we carry out $C_{3}^{1}\times C_{2}^{1}=6$ experiments.
+For SEEG data, we follow the same setting as fNIRS.
+Notice that for SEEG data, inconsistent labels already exist in the raw data.
+We obtain a high-quality testing group by using a majority voting procedure to determine the boundaries.
+Then we leave the testing group aside and only change the validation group to report the mean value of $C_{3}^{2}=3$ experiments.
+All the experimental results are listed in lexicographical order according to the group name composition.
+We also report the mean value and standard derivation of all experiments.
+Specifically, we use the STDEVA to estimate standard deviation based on a sample of data.
 TAS TSC NLL TSC&NLL
@@
 $r\%$ Exp MS- TCN2 AS Former DiffAct Mini Rocket Times Net Patch TST SI- GUA UNI- CON Sel- CL SREA Scale- T Con4m
+0 1 41.39 34.82 41.92 52.89 45.80 30.88 46.79 52.71 47.69 47.40 45.80 52.37
+2 77.19 55.28 53.24 75.76 89.33 52.88 80.21 98.20 93.87 76.48 95.08 95.04
+3 43.71 42.61 45.85 54.11 44.75 42.92 46.89 47.60 46.38 48.14 49.43 48.45
+4 81.92 82.37 65.32 75.00 92.82 55.04 77.44 96.50 81.12 81.04 93.65 98.78
+5 90.67 79.96 70.51 94.37 88.03 70.61 94.05 93.02 87.18 91.70 94.63 94.09
+6 83.89 80.07 63.73 69.91 71.70 59.66 68.27 83.61 81.73 67.06 88.05 93.02
+Avg 69.79 62.52 56.76 70.34 72.07 52.00 68.94 78.61 73.00 68.64 77.77 80.29
+Std 21.56 21.08 11.51 15.47 22.00 13.74 19.01 22.67 20.63 18.01 23.52 23.26
+20 1 42.39 44.86 38.64 47.97 45.96 30.52 47.59 51.25 46.59 45.54 48.25 54.82
+2 68.54 64.63 54.97 71.37 83.16 39.71 81.73 97.72 90.03 70.52 95.35 93.31
+3 46.72 47.95 38.36 49.72 48.64 43.00 51.86 50.40 47.72 53.70 51.06 45.06
+4 87.95 77.25 66.51 69.95 85.91 41.34 76.36 91.20 86.85 73.44 93.33 96.37
+5 82.97 68.53 69.22 86.79 87.36 68.77 88.61 91.35 89.62 85.33 90.48 90.61
+6 71.76 62.32 55.47 54.14 70.13 49.41 64.65 79.70 72.84 67.55 81.81 91.38
+Avg 66.72 60.92 53.86 63.32 70.19 45.46 68.47 76.94 72.28 66.02 76.71 78.59
+Std 18.63 12.38 13.20 15.26 18.77 12.95 16.55 21.05 20.45 14.29 21.48 22.49
+40 1 39.17 42.50 37.82 47.65 44.86 30.87 52.92 48.83 48.60 46.60 50.21 50.78
+2 59.62 59.00 51.47 66.27 73.60 42.93 72.53 92.05 94.60 67.78 94.41 91.36
+3 54.82 46.54 42.00 48.52 43.99 39.89 52.23 49.13 47.87 55.59 48.69 53.33
+4 72.94 78.15 60.46 62.46 87.26 46.10 76.04 75.99 80.98 74.54 93.49 91.57
+5 75.12 75.85 57.81 77.54 79.70 62.93 88.42 93.46 86.63 85.48 87.43 82.62
+6 72.08 62.57 54.23 53.03 71.13 51.42 63.43 79.79 78.15 64.05 81.56 83.48
+Avg 62.29 60.77 50.63 59.25 66.76 45.69 67.60 73.21 72.81 65.67 75.97 75.52
+Std 13.94 14.64 8.95 11.68 18.18 10.87 14.13 19.95 19.85 13.74 21.06 18.58
+Table 5: Full results of the label disturbance experiment on HHAR data. The best results are in bold and we underline the second best results. The worst results are denoted in italics.
+TAS TSC NLL TSC&NLL
+$r\%$ Exp MS- TCN2 AS Former DiffAct Mini Rocket Times Net Patch TST SI- GUA UNI- CON Sel- CL SREA Scale- T Con4m
 0 1 59.46 55.89 51.79 62.16 58.73 58.42 54.79 62.41 63.49 48.95 63.05 68.80
@@
 Table 7: Full results of the label disturbance experiment on SEEG data. The best results are in bold and we underline the second best results. The worst results are denoted in italics.
-Appendix G Full Results
-The full results of the label disturbance experiment are listed in Table 45 and 6.
-For fNIRS, we first divide the data into $4$ groups by subjects and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct cross-validation experiments.
-Therefore, there are $C_{4}^{2}\times C_{2}^{1}=12$ experiments in total.
-Similarly, we divide the Sleep data into $3$ groups and follow the $1$-$1$-$1$ experimental setting.
-Therefore, we carry out $C_{3}^{1}\times C_{2}^{1}=6$ experiments.
-For SEEG data, we follow the same setting as fNIRS.
-Notice that for SEEG data, inconsistent labels already exist in the raw data.
-We obtain a high-quality testing group by using a majority voting procedure to determine the boundaries.
-Then we leave the testing group aside and only change the validation group to report the mean value of $C_{3}^{2}=3$ experiments.
-All the experimental results are listed in lexicographical order according to the group name composition.
-We also report the mean value and standard derivation of all experiments.
-Specifically, we use the STDEVA to estimate standard deviation based on a sample of data.
 Appendix H Case Study